### v1

# SN2023syz and SN2025cbj: Two Type IIn Supernovae Associated with IceCube High-energy Neutrinos

Abstract
Type IIn supernovae (SNe IIn) are a subclass of core-collapse SNe in which strong interactions occur between the ejecta and dense circumstellar material, creating ideal conditions for the production of high-energy neutrinos. This makes them promising candidate sources of neutrinos. In this work, we conduct an association study between 163 SNe IIn observed by the Zwicky Transient Facility and 138 neutrino alert events detected by the IceCube neutrino observatory. After excluding alerts with poor localization, we find two SNe that are spatiotemporally coincident with neutrino events. IC231027A and IC250421A coincide with the positions of SN2023syz and SN2025cbj, respectively, within their localization uncertainties, and the neutrino arrival times are delayed by 38 days and 61 days relative to the discovery times of the corresponding SNe. Using Monte Carlo simulations, we estimate that the probability of such two coincidences occurring by chance in our sample is $p\sim 0.67\%$, suggesting a high likelihood that they arise from genuine associations, though the result is not yet statistical significant. Furthermore, model calculations show that the expected numbers of neutrino events from these SNe IIn could be consistent with the actual observations. Our study provides possible evidence that interacting SNe may be potential neutrino-emitting sources.
Neutrino astronomy
show]liangyf@gxu.edu.cn
show]wangxg@gxu.edu.cn
1 Introduction
[…]
In addition to serving as sources of MeV neutrino emission (as has been shown by Supernova 1987A (K. Hirata et al., 1987)), supernovae have also been extensively studied as potential sources of high-energy (TeV–PeV) neutrinos (V. N. Zirakashvili & V. S. Ptuskin, 2016; K. Murase, 2018; H.-N. He et al., 2018; Z. Li, 2019; P.-W. Chang et al., 2024; R. Abbasi et al., 2023b; S. P. Cosentino et al., 2025). Among all types of supernovae, Type IIn SNe are a subclass of core-collapse supernovae (CCSNe) whose spectra exhibit strong, narrow hydrogen emission lines (“n” standing for “narrow”). These narrow lines, often accompanied by long-lasting and luminous light curves, originate from the collision of high-velocity supernova ejecta with very dense circumstellar material (CSM) shed by the progenitor star shortly before explosion (N. Smith, 2017). Such intense interaction between the supernova ejecta and the dense CSM creates ideal conditions for high-energy neutrino production. As the fast-moving supernova ejecta plows into the dense CSM, powerful shocks can form, accelerating protons to ultra-high energies (K. Murase et al., 2011). These high-energy protons then collide with the dense CSM through proton–proton ($pp$) interactions, producing muons which in turn decay into neutrinos (K. Murase et al., 2011; K. Murase, 2018; A. Kheirandish & K. Murase, 2023; T. Pitik et al., 2023). Because having both an powerful particle accelerator and a dense target, Type IIn supernovae possess the two key ingredients needed for neutrino production, making them stand out among other types of supernovae as promising sources detectable by the IceCube neutrino observatory.
Models show that nearby SNe IIn (tens of Mpc or closer) with efficient CR acceleration could produce a neutrino signal detectable by IceCube (M. Petropoulou et al., 2017). As a population they might contribute a non-negligible fraction of IceCube’s diffuse flux under optimistic assumptions (M. Petropoulou et al., 2017). Single-source and stacking searches conducted with IceCube have yielded meaningful constraints, but so far there is no definitive and widely accepted detection of a Type IIn supernova as a high-energy neutrino source (R. Abbasi et al., 2023b). However, follow-up observations of neutrino alerts do find a few interacting SNe candidates (interacting SNe in general including Ibn/IIn) (M. G. Aartsen et al., 2015b; R. Stein et al., 2025). Recently, R. Stein et al. (2025) found an interacting supernova, SN 2023uqf, in the optical follow-up observations of IceCube alert events with the Zwicky Transient Facility (ZTF), which coincided in time with the high-energy neutrino IC231004A, providing a observational evidence that interacting supernovae can serve as hadronic cosmic-ray accelerators.
In view of the above arguments, a systematic search for associations between SNe IIn and neutrinos is necessary. This will help determine whether SNe IIn are effective high-energy neutrino emitters and quantify their contribution to the IceCube diffuse neutrino flux. In this work, we conduct a systematic search for spatial-temporal coincidences between SNe IIn and IceCube alert events. We use the SNe IIn classified by ZTF - Bright Transient Survey (BTS) (D. A. Perley et al., 2020) as our supernova sample. This is a public catalog with a strict pipeline to conduct a classification for transient sources, and it conducts a magnitude-limited ($m<19\,{\rm mag}$ in either the $g$ or $r$ filter) survey for extragalactic transients in the ZTF public stream (D. A. Perley et al., 2020). We find two spatial-temporal coincidence events, namely IC231027A-SN2023syz ($z$ = 0.037) and IC250421A-SN2025cbj ($z$ = 0.0675). We calculate the model-predicted number of neutrino events for these two sources to check whether they are compatible with observations.
2 Association analysis
2.1 The Neutrino Sample
High-energy astrophysical neutrinos beyond the atmospheric background have been observed by IceCube in both cascade and track data. To help determine the origin of these neutrinos through follow-up observations, since 2016, alerts of individual high-energy neutrino events have been released in real time to platforms for the multi-messenger observation community, such as the General Coordinates Network (GCN) [^1] https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. These alert events mainly focus on track-like neutrino candidates, as they have higher angular resolution than cascade events. In 2019, the alert system was updated. Each alert event is now assigned a “signalness” value (representing the probability that the alert is of astrophysical origin). Based on this value, alerts are categorized as “gold” or “bronze,” corresponding to astrophysical origin probabilities greater than 50% and 30%, respectively (assuming a spectral index of –2.19 for astrophysical neutrinos). For each alert event, information on energy, direction, the 90% position uncertainty ($r_{90}$, statistical error only), and arrival time is reported. For each alert event, updated reports will be issued after further processing provides more accurate information. In our analysis, we use the latest report information for each alert event. We note that one alert event (IC241016222https://gcn.nasa.gov/circulars/37794) was explicitly determined not to be of astrophysical origin, and is therefore excluded from our sample.
We note that R. Abbasi et al. (2023c) has provided alert events from before 2019. However, considering the start time of the ZTF operation, we focus primarily on neutrinos released after 2019, since this period better matches the survey time of the ZTF-BTS catalog. This allows alert events to occur after the SNe listed in the ZTF-BTS catalog. Therefore, for consistency, in this analysis we use only alerts listed in the GCN after 2019. We include all bronze and gold alert events in our sample, which have at least a 30% probability of astrophysical origin. In addition, in the association analysis, good spatial localization of neutrino events is essential, as large localization errors can lead to a high rate of chance coincidences. Many previous studies (e.g., A. Plavin et al. (2020); T. Hovatta et al. (2021)) excluded alert events when the area of their 90% confidence level (C.L.) error region (denoted as $\Omega_{90}$) exceeds a certain threshold. Accordingly, in this analysis, we retain only alert events with $\Omega_{90}\leq 30\,\text{deg}^{2}$ to remove those alerts with large error regions. Our final alert sample contains a total of 138 alert events from June 19, 2019 to June 30, 2025, as shown in Fig. 1.
2.2 The SNe IIn Sample from ZTF-BTS
ZTF is a optical time-domain survey that uses a wide-field (47 deg2) camera on the Samuel Oschin Telescope at Palomar Observatory to systematically scan the night sky (E. C. Bellm et al., 2019). Its primary purpose is to discover transient sources. Since 2018, ZTF has conducted a survey of the visible northern sky ($\sim 3\pi$) with a three-day cadence, and announce newly discovered transient candidates via public alerts (F. J. Masci et al., 2018; M. J. Graham et al., 2019; E. C. Bellm et al., 2019; U. C. Fremling et al., 2020). In December 2020, ZTF entered the second phase (Phase-II) of its public survey operations, with the cadence increased to a two-day cadence. In addition to its photometric survey, ZTF-BTS carries out an extensive spectroscopic campaign aimed at spectroscopically classifying all extragalactic transients with peak magnitudes brighter than 18.5 mag in the $g$ or $r$ bands (and, when spectroscopic resources allow, transients as faint as magnitude 19.0 mag). All classification results are publicly released (U. C. Fremling et al., 2020). The survey has cataloged more than 10,000 objects and is updated nearly daily [^3] https://sites.astro.caltech.edu/ztf/bts/bts.php. The ZTF-BTS catalog adopts a set of strict quality criteria for transient classification, such as magnitude limit, number of observations, and Galactic extinction threshold (see D. A. Perley et al. (2020) for details). Consequently, this source catalog provides a good sample of SNe IIn/candidates with reliable classification for performing the association analysis. Based on the classifications reported in the ZTF-BTS, we select a total of 163 SNe IIn as of June 30, 2025. We consider only sources with high classification reliability exhibiting “SN-like behavior” (slow rise and/or fade time, or coincident with a galaxy). The sky distribution of these SNe is shown in Fig. 1.
Figure 1: Sky distributions in equatorial coordinates of the SNe IIn and alert events adopted in this work. The black points represent the positions of 163 SNe IIn, and the red points represent the locations of 138 alert events.
2.3 Crossmatch and Chance Probability
We perform a systematic search for SNe IIn coincident with neutrino events based on the following criteria: (1) the position of the SN IIn lies within the $r_{90}$ error radius of the neutrino event; (2) the arrival time of the neutrino alert event is within 100 days of the discovery date of the SN IIn. We find that SN 2025cbj and SN 2023syz are spatially and temporally associated with the neutrino events IC250421A and IC231027A, respectively (see Table 1 for their information). IC250421A and IC231027A arrived 61 and 38 days after the discovery date of SN 2025cbj and SN 2023syz (2025 February 19 and 2023 September 19, MJD == 60725 and 60206), respectively. The angular distances between the SNe and the central positions of their corresponding alert events are 1.79 for SN 2025cbj and 2.10 for SN 2023syz.
Table 1: The properties of SN2023syz and SN2025cbj.
$t_{\rm rise,obs}$ Time delay $M_{\rm abs}$ Redshift RA DEC Association $E_{\nu}$
[days] [days] [deg] [deg] [TeV]
SN2023syz 10 38 17.58-17.58 0.037 268.85 45.22 IC231027A 191.5
SN2025cnj 50 61 19.15-19.15 0.0675 239.92 27.11 IC250421A 151.4
We then estimate the chance probability of obtaining two coincidence events using Monte Carlo (MC) simulations. Since IceCube is located at the South Pole and its detection sensitivity depends only on the zenith angle (M. G. Aartsen et al., 2017b, 2020), we generated mock neutrino samples by randomizing the right ascension (RA) while keeping the declination (DEC) and $r_{90}$ error radius fixed at their observed values. The arrival times of the mock neutrinos are uniformly drawn between the earliest and latest arrival times in the dataset. In addition, we generate simulated SN IIn samples by randomizing their positions within a radius of $10^{\circ}$ around their original locations, while requiring that each simulated source remain within the sky coverage of the ZTF-BTS field. We also apply a Kolmogorov–Smirnov (KS) test to the distributions of $\rm RA$ and $\rm DEC$ for the simulated sources to ensure that the mock samples preserve the distribution characteristics of the real sources.
By performing $10^{4}$ Monte Carlo simulations, we estimate the chance probability (i.e., the $p$-value for rejecting the null hypothesis) of finding at least two matches with a time delay of less than 61 days. The probability was computed using p=(M+1)/(N+1)p={(M+1)}/{(N+1)} where $M$ is the number of simulations in which the statistic (i.e., the number of matches) is greater than or equal to the observed value, and $N=10^{4}$ is the total number of MC simulations (A. C. Davison & D. V. Hinkley, 1997). We obtain a chance probability of $p\sim 0.67\%$. This result reflects that the probability of occurring two (or more) coincidence in our sample by chance is low. It is rather likely that the SNe and neutrino associations we have found are physically real, but the current results have not yet reached a level of statistically significant.
3 Model-predicted neutrino numbers from SN 2025cbj and SN 2023syz
The statistical analysis above provides supporting evidence for SNe IIn being sources of high-energy neutrinos. Below, we will investigate whether such a result aligns with model expectations.
We consider a spherically symmetric ejecta propagating into the CSM, with ejecta mass $M_{\rm ej}$ and kinetic energy $E_{k}$. The CSM is assumed to be spherical, steady, and wind-like, with an outer radius $R_{\rm csm}$ and a total mass $M_{\rm csm}$. The number density of the CSM can be expressed as $n_{\rm CSM}(R)={\dot{M}}/{4\pi v_{w}mR^{2}}$, where $\dot{M}$ is the mass-loss rate of the progenitor star, $v_{w}$ is the stellar wind velocity, and $m=\mu m_{\rm p}$. Here $m_{\rm p}$ is the proton mass, $\mu=1.3$ is the mean molecular weight of a neutral gas of solar abundance, and $R$ is the distance from the stellar core. The CSM is shed by the progenitor star before the SN explosion. The stellar wind velocity is taken to be $v_{w}=100\,\rm km\,s^{-1}$ (T. J. Moriya et al., 2014), and the average mass-loss rate is (A. Suzuki et al., 2021)
$\dot{M}=0.3\ {M}_{\odot}\ {\rm yr}^{-1}\left(\frac{M_{\rm CSM}}{10\ M_{\odot}}\right)\left(\frac{R_{\rm CSM}}{10^{16}\ \rm cm}\right)^{-1}\left(\frac{v_{w}}{100\ \rm km\ s^{-1}}\right).$ (1)
The interaction between the ejecta and the CSM creates shocks that accelerate particles. In our calculation, we only consider the forward shock and neglect the contribution from the reverse shock following previous works (e.g., M. Petropoulou et al., 2016; T. Pitik et al., 2022). The temporal evolution of the radius of the shocked shell could be described as (A. Suzuki et al., 2020; R. A. Chevalier & C. Fransson, 2016; T. J. Moriya et al., 2013; T. Pitik et al., 2022)
[…]
where $R_{\rm dec}={M_{\rm ej}u_{w}}/{\dot{M}}$ is the deceleration radius, and $s=10$ is the outer slope of the ejecta density profile. Consequently, the shock velocity is $v_{\rm{sh}}(t)=dR_{\rm sh}/dt\propto t^{(-1)/(s-2)}$ for $R\leq R_{\rm dec}$ and $v_{\rm{sh}}(t)\propto t^{-1/3}$ for $R>R_{\rm dec}$.
During the expansion of the shock, protons are accelerated, and high-energy neutrino emission can be produced through interactions between shock-accelerated protons and cold protons in the dense CSM ($pp$ interaction) (B. Katz et al., 2012; R. Abbasi et al., 2021). At radii smaller than the shock breakout radius $R_{\rm bo}$, where the optical depth satisfies $\tau>c/v_{\rm sh}$, the high-density environment suppresses acceleration. Therefore, efficient particle acceleration only occurs after this radius (T. A. Weaver, 1976; A. Levinson & O. Bromberg, 2008; K. Murase et al., 2011). The shock breakout radius can be determined by solving
$\tau_{T}(R_{\rm{bo}})=\int_{R_{\rm{bo}}}^{R_{\rm{CSM}}}\rho_{\rm{CSM}}(R)\,\kappa_{\rm{es}}\,dR=\frac{c}{v_{\rm{sh}}},$ (3)
where $\kappa_{\rm{es}}\sim 0.34\,\rm{cm}^{2}\rm{g}^{-1}$ (T. Pan et al., 2013) is the electron scattering opacity at solar abundance, $c$ is the speed of light, and $\rho_{\rm CSM}$ is the CSM density at radius $R$.
The evolution of the proton distribution between $R_{\rm bo}$ and $R_{\rm CSM}$ follows (S. J. Sturner et al., 1997; J. D. Finke & C. D. Dermer, 2012; M. Petropoulou et al., 2016):
$\frac{\partial N_{\rm{p}}(\gamma_{\rm{p}},R)}{\partial R}-\dfrac{\partial}{\partial\gamma_{\rm{p}}}\bigg{[}\frac{\gamma_{\rm{p}}}{R}N_{\rm{p}}(\gamma_{\rm{p}},R)\bigg{]}+\frac{N_{\rm{p}}(\gamma_{\rm{p}},R)}{v_{\rm{sh}}(R)\,t_{pp}(R)}=Q_{\rm{p}}(\gamma_{\rm{p}},R),$ (4)
where $N_{\rm{p}}(\gamma_{\rm{p}},R)$ denotes the number density of protons at a given radius $R$ with energies between $\gamma_{\rm{p}}$ and $\gamma_{\rm{p}}+\rm{d}\gamma_{\rm{p}}$. The second term on the left-hand side represents energy losses due to the adiabatic expansion of the SN shell, while $pp$ collisions are treated as an effective escape process, with a characteristic timescale $t_{\rm pp}=(k_{\rm{pp}}\sigma_{\rm{pp}}n_{\rm{sh}}c)^{-1}$ (S. J. Sturner et al., 1997).
For a wind density profile of CSM, the proton injection rate in Eq. (4) is (T. Pitik et al., 2022)
$Q_{\rm p}(\gamma_{\rm p},R)=\frac{9\pi\epsilon_{\rm p}R^{2}_{\rm bo}n_{\rm bo}}{8{\ln}\left({\gamma_{p,\rm{max}}}/{\gamma_{p,\rm{min}}}\right)}\Bigg{[}\frac{v_{\rm sh}(R_{\rm bo})}{c}\Bigg{]}^{2}\Bigg{[}\frac{R}{R_{\rm bo}}\Bigg{]}^{2\alpha}\gamma_{\rm p}^{-k},\quad\text{for}\ \ \gamma_{p,\rm{min}}<\gamma_{p}<\gamma_{p,\rm{max}},$ (5)
where $\alpha$ is the radial dependence index of the shock velocity, $\alpha=-1/7$ (1/2-1/2) for $R<R_{\rm dec}$ ($R>R_{\rm dec}$). $\gamma_{p,\rm{min}}$, $\epsilon_{p}$, $k$ are the minimum Lorentz factor of protons, the proportion of kinetic energy converted into proton acceleration, and the proton spectral index, respectively. We follow the previous works of, e.g., T. Pitik et al. (2022); K. Murase et al. (2011), and set them as ($\gamma_{p,\rm{min}}$, $\epsilon_{p}$, $k$) = (1,0.1,21,0.1,-2). The maximum Lorentz factor of protons $\gamma_{p,\rm max}$ can be calculated by requiring that the acceleration timescale is shorter than the cooling timescale. For proton cooling, we consider the cooling due to inelastic collisions and adiabatic expansion of the shocked shell. More details on the calculation of $\gamma_{p,\rm max}$ can be seen in T. Pitik et al. (2022).
The accelerated protons produce neutrinos through $pp$ interactions. After solving $N_{p}(\gamma_{p},R)$ from Eq. (4), combining it with the cross section of $pp$ interaction $\sigma_{\rm pp}$, the neutrino production rate ($Q_{\nu_{i}+\bar{\nu}_{i}}$ with $i=\mu,e$) can be given by .
$Q_{\nu_{i}+\bar{\nu}_{i}}(E_{\nu},R)=\frac{4n_{\rm{CSM}}(R)m_{\rm{p}}c^{3}}{v_{\rm{sh}}}\int_{0}^{1}{\rm d}(\ln x){\sigma_{\rm{pp}}(E_{\nu}/x)}N_{\rm{p}}\left(\frac{E_{\nu}}{xm_{\rm{p}}c^{2}},R\right)F^{\prime}_{\nu_{i}}(E_{\nu},x)$ (6)
where $x=E_{\nu}/E_{\rm p}$ and $F^{\prime}_{\nu_{i}}(E_{\nu},x)=F^{(1)}_{\nu_{\mu}}(E_{\nu},x)+F^{(2)}_{\nu_{\mu}}(E_{\nu},x)$ for $i=\mu$ and $F^{\prime}_{\nu_{i}}(E_{\nu},x)=F_{\nu_{e}}(E_{\nu},x)$ for $i=e$. The $F^{(1)}_{\nu_{\mu}}$, $F^{(2)}_{\nu_{\mu}}$, and $F_{\nu_{e}}$ adopt the form in S. R. Kelner et al. (2006), which are valid for $E_{p}>0.1$ TeV.
The observed neutrino and antineutrino flux of each flavor ($F_{\nu_{\alpha}+\bar{\nu}_{\alpha}}$ with $\alpha=e,\mu,\tau$) can be calculated as
$F_{\nu_{\alpha}+\bar{\nu}_{\alpha}}(E_{\nu},t)=\frac{(1+z)^{2}}{4\pi d^{2}_{L}(z)}v_{\rm{sh}}(t)\underset{\beta}{\sum}P_{\nu_{\beta}\rightarrow\nu_{\alpha}}Q_{\nu_{\beta}+\bar{\nu}_{\beta}}\left(E_{\nu_{\alpha}}(1+z),R(t)\right),$ (7)
where $P_{\nu_{\beta}\rightarrow\nu_{\alpha}}$ is the flavor change due to neutrino oscillations during the propagation of neutrinos (L. A. Anchordoqui et al., 2014; T. Pitik et al., 2022).
Finally, the expected event rate of muon neutrinos detected by IceCube is
$\dot{N}_{\nu_{\mu}+\bar{\nu}_{\mu}}(t)=\int_{E_{\nu,\rm{min}}}^{E_{\nu,\rm{max}}}dE_{\nu}A_{\rm{eff}}(E_{\nu},\delta)F_{\nu_{\mu}+\bar{\nu}_{\mu}}(E_{\nu},t)$ (8)
where Aeff($E_{\nu}$, $\delta$) is the detector effective area for which we use the effective area given in Ref. (R. Abbasi et al., 2023c). The minimum neutrino energy $E_{\nu,\rm min}$ depends on the effective area, and $E_{\nu,\rm max}$ is related to the maximum protons energy. In Figure 2, we present the neutrino event rate of SN2023syz and SN2025cbj for chosen sets of model parameters. By integrating the temporal interval from the breakout time $t_{\rm{bo}}$ to the arrival dates of alert events $t_{f}$, i.e. $t_{\rm bo}=60206$ (MJD), $t_{f}=60244$ (MJD) for SN2023syz and $t_{\rm bo}=60725$ (MJD), $t_{f}=60786$ (MJD) for SN2025cbj, the total number of neutrinos can be calculated by $N_{\nu_{\mu}+\bar{\nu}_{\mu}}=\int_{t_{\rm{bo}}}^{t_{f}}dt\ \dot{N}_{\nu_{\mu}+\bar{\nu}_{\mu}}(t)$.
Figure 2: Event rates for neutrinos from SN 2023syz (left panel) and SN 2025cbj (right panel) as a function of the time after the shock breakout assuming $E_{k}=10^{52}$ erg and $E_{k}=3\times 10^{52}$ erg, respectively. The blue band denotes the variation arising from varying the $M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$ parameters in the range of ($M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$) $\in$ ($1-20]\,M_{\odot}$, $1-30]\,M_{\odot}$, $1-4]\times 10^{16}\,\rm cm$) for SN2023syz and ($M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$) $\in$ ($_{\odot}, [170]M[1-70]\,M_{\odot}$, $1-4]\times 10^{16}\,\rm cm$) for SN2025cbj. When deriving the blue band, we have excluded the parameters not satisfying $t_{\rm rise}\leq t_{\rm rise,obs}\leq 1.5\times t_{\rm rise}$ and $t_{\rm CSM}-t_{\rm bo}$ greater than the observed time delay, please see the main text and Figure 3 for details.
Figure 3 presents the model-predicted total number of muon (anti)neutrino events detected by IceCube from SN2023syz and SN2025cbj. We choose specific $E_{k}$ and $R_{\rm CSM}$ values and scan the parameter space of ($M_{\rm ej}$, $M_{\rm CSM}$). In the figure, we have excluded those parameters (shaded region) in the parameter space according to the $t_{\rm{rise}}$ (the time interval between the first detection time [^4] Note that using the first detection time is a conservative choice, as emission is usually present before it. and the peak time of the SNe) and $t_{\rm{CSM}}-t_{\rm{bo}}$ (the time delay between $t_{\rm{bo}}$ and $t_{\rm{CSM}}$). According to the light curves of SN2023syz [^5] https://lasair-ztf.lsst.ac.uk/objects/ZTF23abfglcy/ and SN2025cbj [^6] https://lasair-ztf.lsst.ac.uk/objects/ZTF25aagbrpb/, we adopt the $t_{\rm{rise,obs}}$ of 10 days and 50 days for SN2023syz and SN2025cbj, respectively. The $t_{\rm{rise}}$ can be linked to the other quantities through the following relation (S. Ginzburg & S. Balberg, 2012; T. Pitik et al., 2023),
$t_{\rm{rise}}\sim\int_{R{{}_{\rm bo}}}^{R_{\rm ph}}\frac{2(R-R_{\rm bo})k_{\rm T}\rho_{\rm CSM}(R){\rm d}R}{c}$ (9)
where the $R_{\rm{ph}}$ is the photosphere radius at the optical depth $\tau_{\rm{T}}(R_{\rm ph})$ = 1. Then we require the parameter space to satisfy $t_{\rm{rise}}\leq t_{\rm{rise,obs}}\leq 1.5\times t_{\rm{rise}}$ (T. Pitik et al., 2023), namely we allow an error up to 50% on the $t_{\rm{rise}}$ estimated from the model. Furthermore, the $t_{\rm CSM}-t_{\rm bo}$ is required to be larger than the time delay between the arrival dates of alert events and SNe IIn (i.e. $t_{\rm CSM}-t_{\rm bo}\geq$ 38 days for SN2023syz and 61 days for SN2025cbj, respectively). This criterion ensures that the forward shock is still interacting with the CSM at the moment the alert event occurred.
Figure 3: Expected total number of neutrinos from SN 2023syz (left panel) and SN 2025cbj (right panel). We scan the parameters of ($M_{\rm ej}$, $M_{\rm CSM}$) = ($_{\odot}, [120]M[1-20]\,M_{\odot}$; ($M_{\rm ej}$, $M_{\rm CSM}$) = ($_{\odot}, [170]M[1-70]\,M_{\odot}$ ) to show how the expected number relies on these parameters for SN2023syz and SN2025cbj, respectively. The gray shaded region in the figure represents the excluded parameters for which the expected $t_{\rm rise}$ and $t_{\rm{CSM}}-t_{\rm{bo}}$ do not match the observations.
In Figure 3, we can see that the expected number of detected neutrinos depends on the values of $M_{\mathrm{ej}}$ and $M_{\mathrm{CSM}}$. For SN 2023syz, its narrow parameter range is mainly due to the short rise time $t_{\mathrm{rise}}$ of this SN. Within the considered parameter space, using the most optimistic choice of the parameters (i.e., the red part in the parameter space; note that the value of $E_{k}\sim 10^{52}\,\rm erg$ we have adopted for these two sources, especially SN 2023syz, is probably also optimistic), the expected numbers of detected neutrinos can reach $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim$ 0.0016 for SN2023syz and $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim$0.0014 for SN2025cbj. If we simply assume that other SNe IIn also have a comparable neutrino yield (about 0.0015 per source), then for 163 SNe IIn in our sample the total expected number of detected neutrinos is about $\sim 0.24$. This implies that in the optimistic scenario detecting 1–2 neutrino alerts is possible and our model calculation does not rule out the possibility that the two detected neutrino events originated from the SNe.
4 Summary
This study presents a systematic search for spatial-temporal associations between ZTF Type IIn supernovae and high-energy neutrino alerts detected by the IceCube neutrino observatory. SNe IIn are core-collapse supernovae characterized by strong interactions between their ejecta and dense CSM. These interactions create ideal conditions for neutrino production through efficient proton acceleration and $pp$ interactions, making SNe IIn promising candidates for neutrino sources detectable by IceCube. In view of this, we search for spatial and temporal coincidences between 163 SNe IIn observed by ZTF-BTS and 138 IceCube neutrino alerts (2019-2025) with well localization ($\Omega_{90}\leq 30\,\rm deg^{2}$).
Our analysis identify two SNe-neutrino coincidences: SN2023syz ($z=0.037$) with neutrino event IC231027A (38-day delay after supernova onset) and SN2025cbj ($z=0.0675$) with IC250421A (61-day delay). Both neutrinos arrived within the localization uncertainty of their respective supernovae. Monte Carlo simulations estimate that the chance probability of observing $\geq 2$ such coincidences in our sample is only $\sim$0.67%, suggesting these associations may be physical real rather than by chance, though not yet statistically conclusive. To assess the physical plausibility of these associations, we perform model calculations of neutrino emission from these two sources. Using a neutrino production model involving ejecta evolution, shock acceleration, and $pp$ interactions, we compute the expected muon neutrino event rates at IceCube, accounting for CSM properties (mass $M_{\rm CSM}$, radius $R_{\rm CSM}$), ejecta kinetic energy ($E_{k}$) and mass ($M_{\rm ej}$), and observational constraints from the rise time $t_{\rm rise}$ of the optical light curves. We find that, for the most optimistic parameters investigated in this work (the red regions in Figure 3), the predicted numbers of muon neutrino events detectable by IceCube reach $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim 0.0016$ events for SN 2023syz and $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim 0.0014$ events for SN 2025cbj, respectively. This implies that totally detecting 1-2 neutrino alerts from the 163 SNe IIn is possible. The model-expected number of neutrino events from these SNe IIn can accommodate the actual detections.
This work provides supportive (though not conclusive) evidence that interacting supernovae, particularly SNe IIn, could be sources of high-energy neutrinos. Though the limited sample size precludes definitive claims, the relatively low chance probability and model compatibility suggest that SNe IIn may contribute to the high-energy neutrino sky. Future observations with enhanced neutrino sensitivity and more powerful transient surveys (e.g., Vera Rubin Observatory (Ž. Ivezić et al., 2019; K. Bricman & A. Gomboc, 2020)) will help confirm these associations and quantify the role of SNe IIn in cosmic neutrino production.
This work is supported by the National Key Research and Development Program of China (Grant No. 2022YFF0503304), the National Natural Science Foundation of China (12373042, U1938201, 12494573), the Programme of Bagui Scholars Programme (WXG) and Innovation Project of Guangxi Graduate Education (YCBZ2024060).